# Uncertainty-Aware Critic Augmentation for Hierarchical Multi-Agent EV Charging Control

### v4

@@ Abstract
The advanced bidirectional EV charging and discharging technology, aimed at supporting grid stability and emergency operations, has driven a growing interest in workplace applications. It not only reduces electricity expenses but also enhances the resilience in handling practical matters, such as peak power limitation, fluctuating energy prices, and unpredictable EV departures. Considering these factors systematically can benefit energy efficiency in office buildings and for EV users simultaneously. To employ AI to address these issues, we propose HUCA, a novel real-time charging control for regulating energy demands for both the building and EVs. HUCA employs hierarchical actor-critic networks to dynamically reduce electricity costs in buildings, accounting for the needs of EV charging in the dynamic pricing scenario. To tackle the uncertain EV departures, we introduce a new critic augmentation to account for departure uncertainties in evaluating the charging decisions, while maintaining the robustness of the charging control. Experiments on real-world electricity datasets under both simulated certain and uncertain departure scenarios demonstrate that HUCA outperforms baselines in terms of total electricity costs while maintaining competitive performance in fulfilling EV charging requirements. A case study also manifests that HUCA effectively balances energy supply between the building and EVs based on real-time information, showcasing its potential as a key AI-driven solution for vehicle charging control.
keywords: Hierarchical reinforcement learning, EV bidirectional charging, real-time charging control, dynamic pricing
@@ 1 Introduction
Background. The rapid growth in Electric Vehicle (EV) adoption, driven by global sustainability efforts, has led to a surge in demand for advanced charging solutions. According to a report by CBRE Group (Coldwell Banker Richard Ellis Group) Incorporated [4], workplace charging sessions increased at twice the rate of new charging station installations in 2023. This highlights the growing interest for EV charging options in the workplace. Meanwhile, recent advances in bidirectional charging systems [19] [23], which support both Grid-to-Vehicle (G2V) and Vehicle-to-Grid (V2G) power flows, have shown the potential to enhance flexibility with regard to charging stability and emergency situations. The effectiveness of bidirectional charging systems has inspired us to explore how better scheduling of G2V and V2G timings can balance the demand for EV charging and reduce overall electricity costs.
@@
Challenges. For real-time charging control of EVs in various scenarios, previous studies [20] [6] have explored the use of multi-agent reinforcement learning (MARL) techniques to regulate EV charging and discharging actions. However, most existing approaches fail to consider real-world dynamic factors, such as dynamic energy prices and the possibility that EV users may depart earlier than the expected time, which complicate determining optimal control strategies for each EV. Moreover, to avert transformer overloads that could destabilize the power grid [24], it is necessary to impose charging power limits, thereby further complicating the management of EV charging. These dynamics and limitations pose significant challenges in balancing the energy supply between the building and EVs while minimizing electricity costs. It is crucial to recognize that managing charging improperly could result in considerably higher electricity bills, as power companies will levy extra charges due to overconsumption of energy [18].
@@
Proposed Method. To tackle these challenges, we propose HUCA (Hierarchical Multi-Agent Control with Uncertainty-Aware Critic Augmentation), a novel framework designed for real-time charging control. HUCA operates under dynamic pricing and penalty mechanisms, consistent with practical situations. A new hierarchical multi-agent framework is devised to balance the energy supply between the building and EVs in dynamic and uncertain environments. The HUCA framework consists of two levels of control: a high-level agent and multiple low-level agents. The high-level agent determines whether to charge or discharge EVs, which is a collective decision for all EVs. Based on the high-level decision, the low-level agents collaboratively and dynamically modulate the individual charging (or discharging) power level for each EV, ensuring they stay within the specified charging power limits, with the goal of satisfying anticipated demands and avoiding transformer overload.
@@
A key innovation of HUCA lies in its capability to handle EVs that deviate from their expected departure times. These unpredictable departures introduce uncertainty, potentially disrupting low-level agents’ charging decisions. To address this, we propose an uncertainty-aware critic augmentation mechanism, which assesses the likelihood of departure deviation and adjusts the assessments of low-level agents’ charging decisions accordingly. This mechanism achieves two objectives: (i) accounting for uncertainties into the assessment of low-level agents’ decisions, and (ii) limiting the direct effect of uncertainties to the present time slot, ensuring the robustness of future decision assessments. Therefore, HUCA dynamically adapts to arbitrary EV behaviors while maintaining robust charging control. HUCA is designed to minimize the electricity costs of the building while striving to fulfill EV charging demands in a dynamic environment.
@@
  • Real-Time Charging Control: We propose HUCA, a novel hierarchical multi-agent structure that balances energy supply between the building and EVs, which optimizes charging controls on-the-fly.
@@ 2 Preliminaries
This section delineates the essential symbols and their definitions. We also introduce the general formulations of Markov decision process (MDP) and deep deterministic policy gradient (DDPG) related to reinforcement learning techniques.
@@
Let $\mathcal{C}=\{c_{1},c_{2},...,c_{N}\}$ denote a set of charging piles (abbreviated as piles henceforth) located at the office building’s charging station. $\mathcal{V}_{t}=\{v_{1},v_{2},...,v_{M}\}$ specifies a set of EVs docking at piles at time slot $t$, where $M\leq N$. The charging capacity of the charging station is denoted as $\mathcal{P}^{\text{max}}$ (kW), which represents the maximum allowable charging power at the station ($\mathcal{P}^{\text{max}}>0$). In contrast, the maximum allowable discharging power is denoted as $-\mathcal{P}^{\text{max}}$ (kW).
@@
Definition 3. (Charging Power Limitation): To ensure power system stability, the charging power of each charging pile is limited by the maximum charging capacity $\mathcal{P}^{\text{max}}$ of the charging station. Given the number of EVs $|\mathcal{V}_{t}|$ docking at the charging station at time slot $t$, the maximum charging power $P^{\text{pile}}_{t}$ of each pile at time slot $t$ is estimated as $P^{\text{pile}}_{t}=\mathcal{P}^{\text{max}}/|\mathcal{V}_{t}|$. The maximum discharging power of each pile is $-P^{\text{pile}}_{t}$.
To consider the fulfillment of the expected SoC $SoC^{j}_{\text{dep}}$ of each EV $v_{j}$, we estimate the specific charging/discharging power boundary that pile $c_{i}$ provides to the connected EV $v_{j}$ at each time slot $t$, following the settings in  [9, 11]. Specifically, given the minimum and maximum SoC that EV $v_{j}$ can reach at time $t$, denoted by $SoC^{LB}_{t}$ and $SoC^{UB}_{t}$, respectively, the specific charging and discharging power boundaries of charging pile $c_{i}$ are as follows:
$\left\{\begin{matrix}\mathcal{P}^{\max}_{i,t}=\min\bigl{(}P^{\text{pile}}_{t},(SoC^{UB}_{t}-SoC^{j}_{t-1}\cdot C^{i}\cdot\eta^{*})\bigr{)}\\ \mathcal{P}^{\min}_{i,t}=\max\bigl{(}-P^{\text{pile}}_{t},(SoC^{LB}_{t}-SoC^{j}_{t-1}\cdot C^{j}\cdot\eta^{*})\bigr{)}\end{matrix}\right.,$
where $SoC^{j}_{t-1}$ is the SoC of $v_{j}$ at time slot $t-1$, and $C^{j}$ denotes the battery capacity for the EV. $\eta^{*}$ is function of charging efficiency, as defined in  [9, 11].
@@
Decision-making problems are commonly modeled as Markov decision processes (MDPs), which defined by the four-tuple information $(S,A,R,T)$: states $S$, actions $A$, rewards $R$ and the state transition probabilities $T$. The main objective of an MDP is to find a policy that maximizes the cumulative rewards by selecting appropriate actions. The Deep Deterministic Policy Gradient (DDPG) algorithm [13], which is an actor-critic policy-based method that extends stochastic policy gradients to deterministic settings, has shown great effectiveness in complex environments. Specifically, its deep actor network, parameterized by $\theta^{\mu}$, approximates the deterministic policy $\mu_{\theta}:S\to A$. The critic network, parameterized by $\theta^{Q}$, estimates the action-value function $Q(s,a|\theta^{Q})$. Overall, the actor network determines the optimal action for a given state, while the critic network evaluates the action’s quality.
@@
Our work aims to benefit both the office building and EVs by enhancing the original single-agent DDPG framework. We model the problem as a hierarchical MDP to optimize multiple objectives simultaneously, aiming to find the most effective bidirectional charging strategy that reduces building overall electricity costs while considering EV charging needs.
At each time slot $t$, given the charging pile set $\mathcal{C}$, EV set $\mathcal{V}_{t}$, EV charging information $\{\mathcal{I}^{i}|v_{i}\in\mathcal{V}_{t}\}$, and the charging and discharging power boundaries $\{\mathcal{P}^{\max}_{i,t},\mathcal{P}^{\min}_{i,t}|v_{i}\in\mathcal{V}_{t}\}$, the objective is to determine the optimal charging (or discharging) power $\mathcal{P}^{\text{opt}}_{i,t}$ for each EV $v_{i}$ at time slot $t$, where $\mathcal{P}^{\min}_{i,t}\leq\mathcal{P}^{\text{opt}}_{i,t}\leq\mathcal{P}^{\max}_{i,t}$. By determining $\mathcal{P}^{\text{opt}}_{i,t}$, in the certain departure scenario, where the EV’s actual departure time matches the planned one, the goal is to minimize the total electricity cost of the building while ensuring that EV users’ expected SoCs are met at their departure. In scenarios where EVs depart earlier than expected [^1] When the departure occurs later than anticipated, achieving desired SoCs of these EVs is relatively straightforward compared to scenarios of early departure. Consequently, this paper concentrates on addressing the uncertainty of early departure., our objective is to achieve their desired SoCs to the greatest extent while minimizing costs, balancing the trade-off between cost reduction and maximizing EV charging fulfillment.
@@ 3 The HUCA Framework
3.1.1 State representations.
A state $s_{h}$, denoted by $s_{h}=(I^{e}_{t},I^{v}_{t},I^{l}_{t-1})$, describes the status at the time slot $t$. The triplet includes: (i.) the electricity usage of the building and the electricity price, $I^{e}_{t}$; (ii.) EVs situations, $I^{v}_{t}$; and (iii.) the information learned by low-level agents at the previous time slot, $I^{l}_{t-1}$. Details are given in Appendix A.2.
3.1.2 Discrete actions for the high-level agent.
Let $a_{h}\in[0,1]$ denote the action selected by the high-level agent. It is then converted into the discrete high-level action $a^{\text{disc}}_{h}$ to determine whether to charge or discharge EVs, as defined below:
@@
3.1.3 High-level objective and reward function.
The total electricity cost of a building is calculated based on two parts: (i) the total electricity consumption and (ii) the amount of electricity exceeding the contracted capacity [7], which is the upper bound of instant electricity consumption. The objective of the high-level agent is to minimize the total electricity cost.
Let $L_{t}$ represent the electricity load, $\Delta L_{t}$ denote the excess electricity consumption of the building, $p_{t}$ be the electricity price, and $t$ specify the time slot. The reward function $r_{h}(s_{h},a^{\text{disc}}_{h})$, briefed as $r_{h}$, is defined as:
(3.2) $r_{h}=\kappa\cdot\Bigl{(}\underbrace{-\sum_{t^{\prime}=1}^{t}p_{t^{\prime}}\cdot L_{t^{\prime}}-\sum_{t^{\prime}=1}^{t}\max(0,\Delta L_{t^{\prime}})\cdot\varphi}_{\text{electricity cost term}}\Bigr{)}$
@@
(3.3) $\mathcal{L}(\theta^{Q}_{h})=\mathop{\mathbb{E}}_{s_{h},a_{h},r_{h},s^{\prime}_{h}\sim\mathcal{D}_{h}}\biggl{[}\Bigl{(}Q_{h}(s_{h},a_{h})-y_{h}\Bigr{)}^{2}\biggr{]},$
$y_{h}=r_{h}+\gamma\bar{Q}_{h}(s^{\prime}_{h},a^{\prime}_{h})|_{a^{\prime}_{h}=\mu_{h}(s^{\prime}_{h})},$
where $\gamma$ is a discount factor. $\bar{Q}_{h}$ is the target Q-function used to stabilize learning. The replay buffer $\mathcal{D}_{h}$ stores the transition experiences of the high-level agent in the form of tuples $(s_{h},a_{h},r_{h},s^{\prime}_{h})$. Subsequently, the actor network $\theta^{\mu}_{h}$ updates the high-level policy $\mu_{h}$ via gradient descent as follows:
(3.4) $\nabla_{\theta^{\mu}_{h}}J(\mu_{h})\!=\!\!\!\!\!\!\mathop{\mathbb{E}}_{s_{h},a_{h}\sim\mathcal{D}_{h}}\!\nabla_{\theta^{\mu}_{h}}\mu_{h}(a_{h}|s_{h})\nabla_{a_{h}}Q_{h}(s_{h},a_{h})|_{a_{h}=\mu_{h}(s_{h})}.$
@@
Based on the charging or discharging decision made by the high-level action, the goal of the low-level control is to determine the optimal charging (or discharging) power level for each pile, taking into account the uncertainty of EV departures. Each pile $c_{i}\in\mathcal{C}$ is regarded as a low-level agent and its task is formulated as an MDP. A multi-agent structure including multiple piles is deployed for low-level control.
3.2.1 State representations.
For a low-level agent $c_{i}\in\mathcal{C}$, the state $s^{i}_{l}$ observed at time slot $t$ comprises two information $s^{i}_{l}=(I^{v_{j}}_{t},I^{h}_{t})$: (i) the real-time charging information of the connected EV $v_{j}\in\mathcal{V}_{t}$, denoted as $I^{v_{j}}_{t}$; and (ii) the information learned by the high-level agents, denoted as $I^{h}_{t}$. For simplicity of the presentation, we provide additional details in Appendix A.2.
@@
Multi-agent DDPG (MADDPG) [14] extends DDPG into a multi-agent policy gradient algorithm, where decentralized agents learn a centralized critic based on the observations and actions of all agents. Inspired by MADDPG, we formulate the objective of each critic network, $\theta^{Q,i}_{h}$, of a low-level agent $c_{i}\in\mathcal{C}$ by minimizing the following loss:
@@
(3.5) $\mathcal{L}(\theta^{Q,i}_{l})=\mathop{\mathbb{E}}_{\textbf{x},\textbf{a},\textbf{r},\textbf{x}^{\prime}\sim\bigcup_{c_{i}\in\mathcal{C}}\{\mathcal{D}^{i}_{l}\}}\biggl{[}\Bigl{(}Q^{i}_{l}(\textbf{x},a^{1}_{l},...,a^{N}_{l})-y^{i}_{l}\Bigr{)}^{2}\biggr{]},$
$y^{i}_{l}=r^{i}_{l}+\gamma\bar{Q}^{i}_{l}(\textbf{x}^{\prime},a^{\prime 1}_{l},...,a^{\prime N}_{l})|_{a^{\prime j}_{l}=\mu^{j}_{l}(s^{\prime j}_{l})},$
@@
where $\textbf{x}=\{s^{1}_{l},s^{2}_{l},...,s^{N}_{l}\}$ collects the states of all low-level agents. $\mathcal{D}^{i}_{l}$ is the replay buffer of agent $c_{i}$, storing the transition experience with tuples $(s^{i}_{l},a^{i}_{l},r^{i}_{l},s^{{}^{\prime}i}_{l})$. $Q^{i}_{l}$ is the action-value function from the critic network of agent $c_{i}$, and $\bar{Q}^{i}_{l}$ is the target Q-function.

### v5

# Towards Hierarchical Multi-Agent Decision-Making for Uncertainty-Aware EV Charging

@@ Abstract
Recent advances in bidirectional EV charging and discharging systems have spurred interest in workplace applications. However, real-world deployments face various dynamic factors, such as fluctuating electricity prices and uncertain EV departure times, that hinder effective energy management. To address these issues and minimize building electricity costs while meeting EV charging requirements, we design a hierarchical multi-agent structure in which a high-level agent coordinates overall charge or discharge decisions based on real-time pricing, while multiple low-level agents manage individual power level accordingly. For uncertain EV departure times, we propose a novel uncertainty-aware critic augmentation mechanism for low-level agents that improves the evaluation of power-level decisions and ensures robust control under such uncertainty. Building upon these two key designs, we introduce HUCA, a real-time charging control framework that coordinates energy supply among the building and EVs. Experiments on real-world electricity datasets show that HUCA significantly reduces electricity costs and maintains competitive performance in meeting EV charging requirements under both simulated certain and uncertain departure scenarios. The results further highlight the importance of hierarchical control and the proposed critic augmentation under the uncertain departure scenario. A case study illustrates HUCA’s capability to allocate energy between the building and EVs in real time, underscoring its potential for practical use.
keywords: Hierarchical reinforcement learning, Uncertain-aware control, EV bidirectional charging, Real-time charging
@@ 1 Introduction
Background. The rapid growth in Electric Vehicle (EV) adoption, driven by global sustainability efforts, has led to a surge in demand for advanced charging solutions. According to a report by CBRE Group (Coldwell Banker Richard Ellis Group) Incorporated [ev.demandFroChargingPile], workplace charging sessions increased at twice the rate of new charging station installations in 2023. This highlights the growing interest for EV charging options in the workplace. Meanwhile, recent advances in bidirectional charging systems [sami2019bidirectional] [upputuri2023comprehensive], which support both Grid-to-Vehicle (G2V) and Vehicle-to-Grid (V2G) power flows, have shown the potential to enhance flexibility with regard to charging stability and response capabilities. The effectiveness of bidirectional charging systems has inspired us to explore how better scheduling of G2V and V2G timings can balance the demand for EV charging and reduce overall electricity costs.
@@
Challenges. For real-time charging control of EVs in various scenarios, previous studies [saner2022cooperative] [Silva2020CoordinationOE] have explored the use of multi-agent reinforcement learning (MARL) techniques to regulate EV charging actions. However, most existing approaches fail to consider real-world dynamic factors, such as dynamic energy prices and the possibility that EV users may depart earlier than the expected time, which complicate determining optimal control strategies for each EV. Moreover, to avert transformer overloads [^1] Transformer overloads occur when the demand for electricity exceeds the capacity of a transformer, thereby increasing the risk of overheating and subsequent failure. that could destabilize the power grid [visakh2023analysis], it is necessary to impose charging power limits, thereby further complicating the management of EV charging. These dynamics and limitations pose significant challenges in balancing the energy supply between the building and EVs while minimizing electricity costs. It is crucial to recognize that managing charging improperly could result in considerably higher electricity bills, as power companies will levy extra charges due to overconsumption of energy [rosado2020framework].
@@
Proposed Method. To tackle these challenges, we propose HUCA (Hierarchical Multi-Agent Control with Uncertainty-Aware Critic Augmentation), a novel framework designed for real-time charging control. HUCA operates under dynamic pricing and penalty mechanisms, consistent with practical situations. A new hierarchical multi-agent framework is devised to balance the energy supply between the building and EVs in dynamic and uncertain environments. HUCA consists of two levels of control: a high-level agent and multiple low-level agents. The high-level agent determines whether to charge or discharge EVs, which is a collective decision for all EVs. Based on the high-level decision, the low-level agents collaboratively modulate the individual charging (or discharging) power level for each EV, ensuring they stay within the specified charging power limits, with the goal of satisfying anticipated demands and avoiding transformer overload.
@@
A key innovation of HUCA lies in its capability to handle EVs that deviate from their expected departure times. These unpredictable departures introduce uncertainty, potentially disrupting low-level agents’ charging decisions. To address this, we propose an uncertainty-aware critic augmentation mechanism, which assesses the likelihood of departure deviation and adjusts the assessments of low-level agents’ charging decisions accordingly. This mechanism achieves two objectives: (i) accounting for uncertainties into the assessment of low-level agents’ decisions, and (ii) limiting the direct effect of uncertainties to the present time slot, ensuring the robustness of future decision assessments. Therefore, HUCA dynamically adapts to arbitrary EV behaviors while maintaining robust control. HUCA is designed to minimize the electricity costs of the building while striving to fulfill EV charging demands in a dynamic environment.
@@
  • Real-Time Charging Control: We propose HUCA, a novel hierarchical multi-agent structure that balances energy supply between the building and EVs, optimizing charging controls on-the-fly.
@@
2 Related Work
2.1 Real-time Charging Control.
Optimizing charging control considers different pricing mechanisms, such as time-of-use pricing [ting2024online, Zhang2023ElectricVL], real-time pricing [bitencourt2017optimal, selim2021electric], and so on. Also, Day-ahead optimization combined with real-time control has been used to address uncertainties in solar output and EV parking behaviors [Guo2016TwoStageEO], electricity prices [Zhao2017RiskBasedDS] as well as energy demand and PV generation [Yan2019OptimizedOC]. Integrated management systems, additional energy resources or battery energy storage systems have also been developed for the adaptive control [Chen2020PowerCS, Chopra2023CostBA, ChandraMouli2019IntegratedPC].
2.2 Reinforcement Learning for Charging Management.
Single-agent RL, multi-agent RL (MARL), and hierarchical MARL approaches were used to manage charging systems. Single-agent RL represented the control system [Li2020ConstrainedEC] [Mhaisen2020RealTimeSF] under specific considerations, such as ensuring EV battery requirements [Zhang2021CDDPGAD], addressing charger shortage [Li2023OptimalEC], maximizing profits for charging stations [wang2019reinforcement], and minimizing users’ charging costs [Yan2021DeepRL]. MARL modeled fine-grained cooperations [Silva2020CoordinationOE, Li2024MultiAgentGR, Park2022MultiagentDR, Yan2022ACC] among EVs and buildings to tackle safety concerns, power transfer overload [Silva2020CoordinationOE], and users’ charging cost or operating cost [Li2024MultiAgentGR, Park2022MultiagentDR, Yan2022ACC]. Hierarchical MARL further separated system operators and EV users, and coordinated them to minimize demand charges and energy costs [saner2022cooperative].
Although these studies optimized charging control with various aspects, most studies do not address the issue of uncertain EV departures. While some works [Mhaisen2020RealTimeSF, Guo2016TwoStageEO, Park2022MultiagentDR, wang2019reinforcement] consider uncertainties in EV departure, they primarily focused on a single objective, such as maximizing system revenue or minimizing user charging costs.
Furthermore, in many of these studies, uncertainty factors are only encoded as part of the state information or leave them implicit in the stochastic transition dynamics in their RL models, without explicitly modeling or quantifying it, which limits the model’s ability to adaptively adjust charging control under varying levels of departure uncertainty.
In contrast, our work addresses dynamic factors, including fluctuating electricity prices, EV charging demands, and the energy requirements of the office building. More importantly, we explicitly incorporate and quantify EV departure uncertainty in our model, allowing it to directly shape the reinforcement learning decision-making process, while avoiding current uncertainty from misleading future charging decisions. This enables reliable charging control that simultaneously minimizes total electricity costs and accounts for EV users’ charging demands in a dynamic environment.
@@ 3 Preliminaries
We describe key symbols, definitions, and general formulations of the Markov decision process (MDP) and deep deterministic policy gradient (DDPG) in reinforcement learning.
@@
Let $\mathcal{C}=\{c_{1},c_{2},...,c_{N}\}$ denote a set of charging piles (abbreviated as piles henceforth) located at the office building’s charging station. $\mathcal{V}_{t}=\{v_{1},v_{2},...,v_{M}\}$ specifies a set of EVs docking at piles at time slot $t$, where $M\leq N$. The charging capacity of the charging station is denoted as $\mathcal{P}^{\text{max}}$ (kW), which represents the maximum allowable charging power at the station ($\mathcal{P}^{\text{max}}>0$).
@@
Definition 3. (Charging Power Limitation): To ensure power system stability, the charging power of each charging pile is limited by the maximum charging capacity $\mathcal{P}^{\text{max}}$ of the charging station, following the settings in  [Li2023OptimalEC] [li2023constrained]. Given the number of EVs $|\mathcal{V}_{t}|$ docking at the charging station at time slot $t$, the maximum charging power $P^{\text{pile}}_{t}$ of each pile at time slot $t$ is estimated as $P^{\text{pile}}_{t}=\mathcal{P}^{\text{max}}/|\mathcal{V}_{t}|$. The maximum discharging power of each pile is $P^{\text{pile}}_{t}$.
To consider the fulfillment of the expected SoC $SoC^{j}_{\text{dep}}$ of each EV $v_{j}$, we estimate the specific charging/discharging power boundary that pile $c_{i}$ provides to the connected EV $v_{j}$ at each time slot $t$. Specifically, given the minimum and maximum SoC that EV $v_{j}$ can reach at time $t$, denoted by $SoC^{LB}_{t}$ and $SoC^{UB}_{t}$, respectively, the specific charging and discharging power boundaries of pile $c_{i}$ are formulated according to [Li2023OptimalEC][li2023constrained] as:
[…]
where $SoC^{j}_{t-1}$ is the SoC of $v_{j}$ at time t1t-1, and $C^{j}$ denotes the battery capacity for $v_{j}$. $\eta^{*}$ is function of charging efficiency.
@@
Decision-making problems are commonly modeled as MDPs, defined by the four-tuple information $(S,A,R,T)$: states $S$, actions $A$, rewards $R$, and the state transition probabilities $T$. The main objective of an MDP is to find a policy that maximizes the cumulative rewards by selecting appropriate actions. The Deep Deterministic Policy Gradient (DDPG) algorithm [lillicrap2015continuous], which is an actor-critic policy-based method that extends stochastic policy gradients to deterministic settings, has shown great effectiveness in complex environments. Specifically, its deep actor network, parameterized by $\theta^{\mu}$, approximates the deterministic policy $\mu_{\theta}:S\to A$. The critic network, parameterized by $\theta^{Q}$, estimates the action-value function $Q(s,a|\theta^{Q})$. Overall, the actor network determines the optimal action for a given state, while the critic network evaluates the action’s quality.
@@
Our work aims to benefit both the office building and EVs by enhancing the original single-agent DDPG framework. We model the problem as a hierarchical MDP to optimize multiple objectives simultaneously, aiming to find the most effective bidirectional charging strategy that reduces building overall electricity costs while considering EV charging needs. At each time slot $t$, given the charging pile set $\mathcal{C}$, EV set $\mathcal{V}_{t}$, EV charging information $\{\mathcal{I}^{j}|v_{j}\in\mathcal{V}_{t}\}$, and the charging and discharging power boundaries $\{\mathcal{P}^{\max}_{i,t},\mathcal{P}^{\min}_{i,t}|c_{i}\in\mathcal{C}\}$, the objective is to determine the optimal charging (or discharging) power $\mathcal{P}^{\text{opt}}_{i,t}$ for the EV connected to pile $c_{i}$ at time slot $t$, where $\mathcal{P}^{\min}_{i,t}\leq\mathcal{P}^{\text{opt}}_{i,t}\leq\mathcal{P}^{\max}_{i,t}$.
By determining $\mathcal{P}^{\text{opt}}_{i,t}$, in the certain departure scenario, where the EV’s actual departure time matches the planned one, the goal is to minimize the total electricity cost of the building while ensuring that EV users’ expected SoCs are met at their departure. In uncertain departure scenario, where EVs depart earlier than expected [^2] When the departure occurs later than anticipated, achieving desired SoCs of these EVs is relatively straightforward compared to scenarios of early departure. Consequently, this paper concentrates on addressing the uncertainty of early departure., our objective is to achieve their desired SoCs to the greatest extent while minimizing costs, balancing the trade-off between cost reduction and maximizing EV charging fulfillment.
@@ 4 The HUCA Framework
State representations. A state $s_{h}$ represents the status at time slot $t$, consisting of three main information $s_{h}=(I^{e}_{t},I^{v}_{t},I^{l}_{t-1})$:
  • $I^{l}_{t-1}$ includes the average and the standard deviation of critic value of all low-level agents from at time t1t-1.
Discrete actions for the high-level agent. Let $a_{h}\in[0,1]$ denote the action selected by the high-level agent. It is then converted into the discrete high-level action $a^{\text{disc}}_{h}$ to determine whether to charge or discharge EVs, as defined below:
@@
High-level objective and reward function. The total electricity cost of a building is calculated based on two parts: (i) the total electricity consumption and (ii) the amount of electricity exceeding the contracted capacity [^3] Typically, the electricity provider establishes a maximum power limit for each subscriber to efficiently manage overall consumption. If the instant electricity consumption of a subscriber surpasses this agreed-upon capacity, they will incur additional penalties. [Fernndez2013CostOO]. The objective of the high-level agent is to minimize the total electricity cost.
Let $L_{t}$ represent the electricity load, $\Delta L_{t}$ denote the excess electricity consumption of the building, $p_{t}$ be the electricity price, and $t$ specify the time slot. Define $L_{t^{\prime}}]_{+}=\max(0,\Delta L_{t^{\prime}})$. The reward function, $r_{h}(s_{h},a^{\text{disc}}_{h})$, abbreviated as $r_{h}$, is formulated as:
(4.2) $\hskip-2.84544ptr_{h}=\kappa\cdot\Bigl(\underbrace{-\sum_{t^{\prime}=1}^{t}p_{t^{\prime}}\cdot L_{t^{\prime}}-[L_{t^{\prime}}]_{+}\cdot\varphi}_{\text{electricity cost term}}\Bigr)$
@@
(4.3) $\hskip-2.84544pt\mathcal{L}(\theta^{Q}_{h})=\mathop{\mathbb{E}}_{s_{h},a_{h},r_{h},s^{\prime}_{h}\sim\mathcal{D}_{h}}\biggl[\Bigl(Q_{h}(s_{h},a_{h})-y_{h}\Bigr)^{2}\biggr],$
$\hskip-2.84544pty_{h}=r_{h}+\gamma\bar{Q}_{h}(s^{\prime}_{h},a^{\prime}_{h})|_{a^{\prime}_{h}=\bar{\mu}_{h}(s^{\prime}_{h})},$
where $\gamma$ is a discount factor. $s^{\prime}_{h}$ is the next state after taking action $a_{h}$. $\bar{Q}_{h}$ and $\bar{\mu}_{h}$ are the target action-value function and target policy, respectively, used to stabilize training. The replay buffer $\mathcal{D}_{h}$ stores the transition experiences of the high-level agent in the form of tuples $(s_{h},a_{h},r_{h},s^{\prime}_{h})$. Subsequently, the actor network, with parameters $\theta^{\mu}_{h}$, refines the high-level policy $\mu_{h}$ via gradient descent as follows:
(4.4) $\nabla_{\theta^{\mu}_{h}}J(\mu_{h})\!=\!\!\!\!\!\!\mathop{\mathbb{E}}_{s_{h}\sim\mathcal{D}_{h}}\!\nabla_{\theta^{\mu}_{h}}\mu_{h}(a_{h}|s_{h})\nabla_{a_{h}}Q_{h}(s_{h},a_{h})|_{a_{h}=\mu_{h}(s_{h})}.$
Using mini-batch updates, the high-level agent’s training loss is computed over transition experiences drawn from the replay buffer $\mathcal{D}_{h}$. At each training step, given a mini-batch $\mathcal{B}_{h}$ of transitions $\bigl\{(s_{h}^{(b)},a_{h}^{(b)},r_{h}^{(b)},s_{h}^{(b)})\bigr\}_{b=1}^{|\mathcal{B}_{h}|}$ drawn from $\mathcal{D}_{h}$, the critic and actor networks, parameterized by $\theta^{Q}_{h}$ and $\theta^{\mu}_{h}$, are updated as shown in Algorithm 1.
Algorithm 1 High-Level Agent Update
1:procedure UpdateHigh($\mathcal{B}_{h}$, $\theta^{Q}_{h}$, $\theta^{\mu}_{h}$)
2:  Update the critic network by minimizing: $\mathcal{L}(\theta^{Q}_{h})=\frac{1}{\mathcal{B}_{h}}\sum_{b}\left(Q_{h}(s^{(b)}_{h},a^{(b)}_{h})-y{(b)}_{h}\right)^{2}$
3:  Update the actor netwok using: $\nabla_{\theta^{\mu}_{h}}J\approx\frac{1}{\mathcal{B}_{h}}\sum_{b}\nabla_{\theta^{\mu}_{h}}\mu_{h}(s^{(b)}_{h})\nabla_{a_{h}}Q_{h}(s^{(b)}_{h},a^{(b)}_{h})$
4:end procedure
@@
Based on the charging or discharging decision made by the high-level action, the goal of the low-level control is to determine the optimal charging (or discharging) power level for each pile, taking into account the uncertainty of EV departures. Each pile $c_{i}\in\mathcal{C}$ is considered as a low-level agent and its task is formulated as an MDP. A multi-agent structure including multiple piles is deployed for low-level control.
State representations. For a low-level agent $c_{i}\in\mathcal{C}$, the state $s^{i}_{l}$ observed at time slot $t$ can be described as comprising two main information $s^{i}_{l}=(I^{v_{j}}_{t},I^{h}_{t})$. Details are:
@@
Multi-agent DDPG (MADDPG) [lowe2017multi] extends DDPG into a multi-agent policy gradient algorithm, where decentralized agents learn a centralized critic based on the observations and actions of all agents. Inspired by MADDPG, we formulate the objective of each critic network, $\theta^{Q,i}_{h}$, of a low-level agent $c_{i}\in\mathcal{C}$ by minimizing the following loss:
@@
(4.5) $\hskip-2.84544pt\mathcal{L}(\theta^{Q,i}_{l})=\mathop{\mathbb{E}}_{\textbf{x},\textbf{a},\textbf{r},\textbf{x}^{\prime}\sim\mathcal{D}_{l}}\biggl[\Bigl(Q^{i}_{l}(\textbf{x},a^{1}_{l},.,a^{N}_{l})-y^{i}_{l}\Bigr)^{2}\biggr],$
$\hskip-2.84544pty^{i}_{l}=r^{i}_{l}+\gamma\bar{Q}^{i}_{l}(\textbf{x}^{\prime},a^{\prime 1}_{l},.,a^{\prime N}_{l})|_{a^{\prime j}_{l}=\bar{\mu}^{j}_{l}(s^{\prime j}_{l})},$
@@
where $\textbf{x}=\{s^{1}_{l},s^{2}_{l},...,s^{N}_{l}\}$ collects the states of all low-level agents, and $\textbf{x}^{\prime}$ contains all the next states. $\mathcal{D}_{l}$ is the replay buffer shared between low-level agents, storing the transition experience with tuples $(\textbf{x},a^{1}_{l},...,a^{N}_{l},r^{1}_{l},...,r^{N}_{l},\textbf{x}^{\prime})$. $Q^{i}_{l}$ is the action-value function from the critic network of agent $c_{i}$. $\bar{Q}^{i}_{l}$ and $\bar{\mu}^{j}_{l}$ are the target action-value function and the target policy, respectively.
Figure 3: The illustration of uncertainty-aware critic augmentation. For each low-level agent $c_{i}$, the action-value function is augmented based on the uncertainty factor (Eq. 4.6). This augmentation guides the updated policy to make decisions with explicit consideration of EV departure uncertainty.